OpenStudy (anonymous):

I need help finding an equation and writing it in both slope intercept form and standard form. Through (6,-3) and (8,4)

OpenStudy (campbell_st):

find the gradient \[m = \frac{ y _{2} -y _{1}}{x _{2} - x}\] \[m =\frac{( 4 - -3)}{(8-6)}\] m = 7/2 then the equation of the line is \[y =\frac{7}{2} x + b\] substitute x = 8 and y = 4 from the 2nd point into the equation above and solve for b. this will give the equation of the line in slope- intercept form

OpenStudy (anonymous):

i got 7/2x+24 for slope intercept

OpenStudy (anonymous):

I got 7x+2y=64 for the standard form but I wasn't sure if it was correct..

OpenStudy (campbell_st):

isn't it 4 = 7/2 * 8 + b 4 = 28 + b subtract 28 from both sides so -24 = b then the equation is y = 7/2 x - 24

OpenStudy (anonymous):

Yes you're right ..small mistake

OpenStudy (campbell_st):

thats ok... happens to us all

OpenStudy (anonymous):

Is the standard form 7x+2y=64

OpenStudy (campbell_st):

y = 7/2 x - 24 multiply every term by 2 2y = 7x - 48 add 48 to both sides and subtract 2y from both sides 7x - 2y = 48

OpenStudy (anonymous):

doesn't it have to be in Ax+By=C form so it would be 7x+2y=48

OpenStudy (campbell_st):

you need to rearrange the equation... in my solution A = 7, B = -2 and C = 48

OpenStudy (campbell_st):

thats just the general form, A,B,C can be positive or negative

OpenStudy (anonymous):

can you help me with one more problem?